Carmelita Inc., has the following information available:

Costs from
Beginning Inventory
Costs from
Current Period
Direct materials $2,000          $22,252     
Conversion costs 6,200          150,536     

At the beginning of the period, there were 500 units in process that were 60% complete as to conversion costs and 100% complete as to direct materials costs. During the period, 4,500 units were started and completed. Ending inventory contained 340 units that were 30% complete as to conversion costs and 100% complete as to materials costs. The company uses the FIFO process cost method.

The cost of completing a unit during the current period was

a.$34.88

b.$36.19

c.$35.95

d.$35.89
Answer #1

Equivalent unit of material = 0+4500+340 = 4840

Equivalent unit of conversion = 500*40%+4500+340*30% = 4802

Cost per equivalent unit of material = 22252/4840 = 4.60

Cost per equivalent unit of conversion = 150536/4802 = 31.35

Cost of completing a unit = 4.60+31.35 = 35.95

So answer is c) $35.95